OpenStudy (anonymous):

what is the x intercept of y= –16x^2 + 210x

hero (hero):

Set y = 0 and solve for x

OpenStudy (anonymous):

i dont no how can you explain ?

hero (hero):

0 = -16x^2 + 210x Solve that

OpenStudy (tasreeb):

y=0, -16x^2+210x x(-16x+210) x=-210/-16

OpenStudy (anonymous):

i got x=0, 105/8

OpenStudy (anonymous):

im assiming im wrong and @tasreeb is right

hero (hero):

0 = -16x^2 + 210x 0 = -x(16x - 210) 0 = x(16x - 210) x = 0 16x - 210 = 0 16x = 210 x = 210/16 x = (0, 210/16)

OpenStudy (anonymous):

so he was wrong also because his answers are negative, ?

hero (hero):

-210/-16 = 210/16

hero (hero):

-/- = +

OpenStudy (anonymous):

oh ya cause two negaties equal a positive lol sorry brain stub !

OpenStudy (tasreeb):

x=0, x=210/16